# Adjust for Trust: Mitigating Trust-Induced Inappropriate Reliance on AI Assistance

### v1

@@ Abstract
Trust biases how users rely on AI recommendations in AI-assisted decision-making tasks, with low and high levels of trust resulting in increased under- and over-reliance, respectively. We propose that AI assistants should adapt their behavior through trust-adaptive interventions to mitigate such inappropriate reliance. For instance, when user trust is low, providing an explanation can elicit more careful consideration of the assistant’s advice by the user. In two decision-making scenarios—laypeople answering science questions and doctors making medical diagnoses—we find that providing supporting and counter-explanations during moments of low and high trust, respectively, yields up to 38% reduction in inappropriate reliance and 20% improvement in decision accuracy. We are similarly able to reduce over-reliance by adaptively inserting forced pauses to promote deliberation. Our results highlight how AI adaptation to user trust facilitates appropriate reliance, presenting exciting avenues for improving human-AI collaboration.
Adjust for Trust:
Mitigating Trust-Induced Inappropriate Reliance on AI Assistance
Tejas Srinivasan  and Jesse Thomason University of Southern California {tejas.srinivasan,jessetho}@usc.edu
1 Introduction
@@
AI systems are being deployed to assist humans in a wide range of decision-making tasks Cai et al. (2019); Chiang et al. (2023); Che et al. (2024). AI-assisted decision-making Lai et al. (2023) typically consists of an AI system providing a recommendation for the user’s consideration. A key factor modulating how users incorporate AI advice is user trust, which is the user’s belief that the AI will help them achieve their goals in situations of uncertainty and vulnerability Lee and See (2004). Having higher trust makes users more likely to accept the AI’s recommendation, all else being equal Dzindolet et al. (2003). Moreover, trust is not a static belief, but instead continuously evolves as the user interacts with the AI and observes decision outcomes Dhuliawala et al. (2023).
@@
Figure 1: User trust in AI systems evolves over a series of decision-making interactions, impacting how carefully the user considers future AI recommendations. To mitigate the effects of extreme trust and encourage critical reliance, AI systems should adapt their behavior to users’ trust levels. For instance, when trust is low, providing explanations reduces under-reliance.
@@
User trust does not always align with AI assistant trustworthiness, i.e. its true capability to help the user Wright (2010). Miscalibrated trust Jacovi et al. (2021) may develop due to recency bias, the user’s internal biases towards AI, or the assistant’s inability to communicate its reasoning or limitations. Miscalibrated trust acts as a cognitive bias Lee (2024) and hinders critical evaluation of AI recommendations, resulting in inappropriate reliance Parasuraman and Riley (1997). For example, we find that doctors mistakenly accept 26% of AI misdiagnoses when their trust is high, compared to 8% when trust is lower, indicating over-reliance. Conversely, when user trust is low, doctors reject correct AI diagnoses 68% of the time, up from 40% otherwise, indicating a bias towards under-reliance.
@@
We posit that AI assistants should adapt their behavior in response to users’ trust levels in order to mitigate inappropriate reliance caused by extreme (low or high) trust. For instance, when trust is low, the assistant can reduce risk of disuse by providing the user with additional reasoning to support its recommendation (Figure 1). Similarly, when users are too trusting, the assistant can highlight reasons its recommendation may be incorrect, or can simply slow down the interaction. We hypothesize that strategically introducing these trust-adaptive interventions will prompt users to engage more carefully with AI advice, rather than accepting or rejecting advice without due consideration.
We examine the effect of trust-adaptive AI interventions at mitigating inappropriate reliance on two decision-making tasks: answering science trivia questions and making medical diagnoses based on patient symptoms. We first validate our premise that, when interacting with the AI assistant over a sequence of decision-making problems, users’ trust level at the start of a given interaction affects their reliance behavior and decision-making performance (§3.2), with extreme levels of user trust (too high or too low) resulting in increased inappropriate reliance. Through controlled studies, we find that strategically providing supporting explanations when user trust is low reduces under-reliance and improves decision-making accuracy (§4.1). Similarly, providing counter-explanations reduces over-reliance when trust is high (§4.2). Combining these interventions to mitigate under- and over-reliance yields complementary improvements in decision-making accuracy and inappropriate reliance (§4.3). We also evaluate the utility of intervening by decelerating the interaction, finding that it helps reduce over-reliance but not under-reliance (§4.4).
@@
Our findings highlight the utility of adapting to user trust in AI-assisted decision-making and present exciting avenues for facilitating appropriate reliance and improving human-AI collaboration.
2 Related Work
Figure 2: In our user study, each user interacts with an AI for a sequence of $30$ decision-making problems. In each problem, the user first makes a decision by themselves, and then receives advice from the AI which they use to make a final decision. The user is then told what the correct decision is, and reports their trust in the AI (out of 10).
@@ 2 Related Work
We draw on a vast literature on measuring user trust in AI systems and evaluate how decision aids can be used for mitigating inappropriate reliance in moments of low or high trust.
@@
Much work has explored the nature of human trust in AI systems Lai et al. (2023), particularly in situations characterized by risk and uncertainty Jacovi et al. (2021). Lee and See (2004) provide the most commonly accepted definition of trust, as a person’s attitude that an agent will help them achieve their goals. User trust is considered to be calibrated  Alizadeh et al. (2022) when it aligns with the AI system’s true capabilities Wright (2010), thus reducing AI misuse and disuse Alizadeh et al. (2022). While trust in AI has typically been attributed to socio-economic and individual factors Bach et al. (2024), recent work Dhuliawala et al. (2023); Pareek et al. (2024) examines how trust develops as users interact with AI systems over multiple timesteps.
@@
Bach et al. (2024) identify a variety of mechanisms for measuring user trust, such as questionnaires Schaffer et al. (2019), qualitative interviews Barda et al. (2020), surveys Lin et al. (2019), and point scales Gulati et al. (2019). In AI-assisted decision-making, early works measured trust by observing user reliance behavior Yin et al. (2019); Zhang et al. (2020); however, de Fine Licht and Brülde (2021) distinguish reliance, an observable behavior, from trust, a subjective belief. Self-reported trust levels from users, where users report their confidence in the AI’s accuracy for a question, are a reasonable proxy for trust, albeit at a local, interaction level Pareek et al. (2024); Dhuliawala et al. (2023). Instead, we adopt a more global lens for eliciting trust scores by asking users to report their belief in the AI’s helpfulness on a scale of 0 to 10.
@@
Inappropriate reliance, where users mistakenly accept incorrect AI predictions or reject correct ones Parasuraman and Riley (1997), is highly undesirable in high-stakes domains, such as healthcare and law Schemmer et al. (2023). Appropriate reliance can be fostered through various decision aids, such as model confidences Zhang et al. (2020); Vodrahalli et al. (2022), explanations Wang and Yin (2021); Bansal et al. (2021), uncertainty expressions Zhou et al. (2024); Kim et al. (2024), and providing sources Feng and Boyd-Graber (2019). Cognitive forcing functions Buçinca et al. (2021), which insert friction Chen and Schmidt (2024); İnan et al. (2025) and promote deliberation Park et al. (2019); Rastogi et al. (2022); Ma et al. (2024a), are effective at mitigating over-reliance. We demonstrate how strategically providing these decision aids to users during moments of low or high trust can mitigate trust-induced inappropriate reliance.
3 How does Trust Impact Reliance on AI?
We study how user trust impacts reliance on AI advice over a sequence of decision-making problems (§3.1). User studies reveal that trust being too high or too low increases inappropriate reliance (§3.2).
3.1 Sequential AI-Assisted Decision-Making
We consider a setting where a human user interacts with an AI assistant on a sequence of $N$ decision-making problems, where each problem belongs to the same task, such as making medical diagnoses based on symptoms. Problems are tuples of input $x$, categorical choices $\mathcal{Y}$, and correct choice $y^{*}\in\mathcal{Y}$. The user solves each problem in three stages: independent decision-making based on their own knowledge, decision revision after viewing the AI recommendation, and trust update in the AI after observing decision accuracy (Figure 2). [^1] Our setup assumes that both the user and AI can observe the ground-truth decision after each problem.
@@ 1. Independent Decision-Making:
For the $i^{th}$ decision-making problem with input $x_{i}$, the user initially makes a decision $y^{u,\text{init}}_{i}\in\mathcal{Y}$.
@@ 2. AI-Informed Decision Revision:
The user then views the AI prediction $y^{AI}_{i}$ and confidence estimate $c^{AI}_{i}$, and makes a final decision $y^{u,\text{fin}}_{i}$.
@@ 3. Trust Update:
After the user makes their final decision $y^{u,\text{fin}}_{i}$, the interface informs the user of the accuracy of their decision and the AI prediction $y^{AI}_{i}$. Observing this feedback may alter the user’s trust in the AI’s ability to help them make better decisions. For instance, the AI misleading the user into making a wrong decision is likely to decay trust. After showing this feedback, we ask users to report how much they trust the AI to help them with the decision-making task based on all user-AI interactions so far, as an integer between 0 and 10.
@@
Our trust operationalization captures a global belief in the AI’s helpfulness, which is likely to influence how the user relies on AI advice in subsequent decision-making problems.
Evaluating Appropriate Reliance.
We only evaluate interactions where the user’s initial decision differs from the AI prediction, i.e. $y^{u,\text{init}}_{i}\neq y^{AI}_{i}$. We capture the degree of users’ reliance on AI assistance using Switch Rate Yin et al. (2019); Zhang et al. (2020), the fraction of interactions where the user switched their decision to the AI prediction. Following Ma et al. (2024b), we capture the degree of appropriate reliance on AI assistance using two metrics: Over-Reliance represents the fraction of interactions where the user switches to the AI’s prediction when the AI is incorrect, while Under-Reliance represents the fraction of interactions where the user does not switch to the AI’s prediction when the AI was in fact correct.
We hypothesize that very high and very low values of trust will increase the probability of inappropriate reliance in the next user-AI interaction. Specifically, we hypothesize that low user trust biases users towards rejecting correct AI advice, i.e. higher Under-Reliance. Similarly, high trust leads to users accepting incorrect advice more frequently, i.e. higher Over-Reliance.
3.2 Experiments
We evaluate the impact of users’ trust level [^2] Henceforth, “user trust” refers to the user’s trust level at the start of an interaction, i.e. the trust score reported by the user at the end of the previous interaction. on their reliance behavior and decision-making performance in subsequent user-AI interactions.
Decision-Making Tasks.
We conduct user studies on two decision-making tasks. The ARC task consists of answering science questions from the ARC dataset Clark et al. (2018). Each problem consists of a question and two options. The Diagnosis task involves making medical diagnoses based on patient intake forms, sourced from the DDXPlus dataset Fansi Tchango et al. (2022). Users must select from four possible diagnoses. Appendix A contains more details about problem selection.
Figure 3: Calibration curves and Expected Calibration Error (ECE) of our simulated AI assistants.
Simulated AI.
@@ 3.2 Experiments
Our user studies use a simulated AI that provides a recommendation $R_{i}=(y^{AI}_{i},c^{AI}_{i})$ for each decision-making problem. We experiment with two types of AI assistants: one perfectly calibrated and one overconfident.
@@
This sampling procedure generates AI predictions and confidence scores for each decision-making problem. Figure 3 shows calibration curves for the calibrated and overconfident AI predictions.
Experiment Setup.
We perform user studies on the Prolific platform, on three task settings: the ARC task with a calibrated AI (ArcC), the ARC task with an overconfident AI (ArcO), and the Diagnosis task with a calibrated AI (DiagC).

### v2

@@ Abstract.
User trust biases reliance on AI assistance during decision-making tasks, with overly low and high levels of trust resulting in increased under- and over-reliance, respectively. We propose that AI assistants should adapt their behavior through trust-adaptive interventions to mitigate such inappropriate reliance. For instance, when user trust is low, that providing an AI explanation can elicit more careful consideration of the assistant’s advice by the user. In two decision-making scenarios—laypeople answering science questions and medical doctors making diagnoses—we find that providing supporting and counter-explanations during moments of low and high trust, respectively, yields up to 38% reduction in inappropriate reliance and 20% improvement in decision accuracy. We are similarly able to reduce over-reliance by adaptively inserting forced pauses to promote deliberation. Our results highlight how AI adaptation to user trust facilitates appropriate reliance, presenting exciting avenues for improving human-AI collaboration.
AI-assisted decision-making, appropriate reliance, user trust
journalyear: 2026copyright: ccconference: 31st International Conference on Intelligent User Interfaces; March 23–26, 2026; Paphos, Cyprusbooktitle: 31st International Conference on Intelligent User Interfaces (IUI ’26), March 23–26, 2026, Paphos, Cyprusdoi: 10.1145/3742413.3789136isbn: 979-8-4007-1984-4/2026/03ccs: Human-centered computing Empirical studies in HCI
1. Introduction
@@
AI systems are being deployed to assist humans in a wide range of decision-making tasks (Cai et al., 2019; Chiang et al., 2023; Che et al., 2024). AI-assisted decision-making (Lai et al., 2023) typically consists of an AI system providing a recommendation for a human user’s consideration. A key factor modulating how users incorporate AI advice is user trust, which is the user’s belief that the AI will help them achieve their goals in situations characterized by uncertainty and vulnerability (Lee and See, 2004). Having higher trust makes users more likely to accept the AI’s recommendation, all else being equal (Dzindolet et al., 2003). Trust is not a static belief; it evolves as the user interacts with the AI and observes decision outcomes (Dhuliawala et al., 2023).
@@
Figure 1. User trust in AI systems evolves over a series of decision-making interactions, impacting how carefully the user considers future AI recommendations. To mitigate the effects of extreme trust and encourage critical reliance, AI systems should adapt their behavior to users’ trust levels.
Evolution of user trust in AI assistants, and how AI assistants can adapt their behavior in response to user trust levels.
@@
User trust does not always align with the AI assistant’s trustworthiness, i.e. its true capability to help the user (Wright, 2010). Miscalibrated trust (Jacovi et al., 2021) may develop due to recency bias, the user’s internal biases towards AI, or the assistant’s inability to communicate its reasoning or limitations. Miscalibrated trust acts as a cognitive bias (Lee, 2024) that hinders critical evaluation of AI recommendations, resulting in inappropriate reliance (Parasuraman and Riley, 1997). For example, we find that medical doctors mistakenly accept 26% of AI misdiagnoses when their trust is high, compared to 8% when trust is lower, indicating over-reliance. Conversely, when user trust is low, doctors reject correct AI diagnoses 68% of the time, up from 40% otherwise, indicating a bias towards under-reliance.
@@
We posit that AI assistants should adapt their behavior in response to users’ trust levels in order to mitigate inappropriate reliance caused by extreme (low or high) trust. For instance, when trust is low, the assistant can reduce risk of disuse by providing the user with additional reasoning to support its recommendation (Figure 1). Similarly, when users are too trusting, the assistant can highlight reasons its recommendation may be incorrect via counter-explanations (Si et al., 2024), or can simply slow down the interaction. We hypothesize that strategically deploying trust-adaptive interventions will prompt users to engage more carefully with AI advice and yield higher decision-making accuracy.
In this work, we conduct user studies on two decision-making tasks: laypeople answering science questions and medical doctors making clinical diagnoses based on patient symptoms. Through controlled studies using a simulated AI assistant, we first validate our premise that user trust levels affect their reliance behavior (Section 5). Our analysis reveals that users’ decisions to rely on AI assistance are indeed strongly correlated with their trust level at the start of an interaction. These findings highlight that trust is an important factor modulating user reliance on AI assistance, and that extreme levels of user trust result in inappropriate reliance of different kinds.
The following experiments (Sections 6, 7, 8, and 9) investigate the utility of trust-adaptive interventions at mitigating inappropriate reliance caused by extreme levels of user trust. We find that adapting AI behavior during at moments of low or high trust by surfacing different kinds of explanations or slowing down interactions is effective at promoting careful deliberation by users and reducing inappropriate reliance. Our findings generalize to a setting where users collaborate with a Large Language Model (LLM) assistant (Section 10), indicating that designing LLMs to be trust-aware presents a promising avenue for improving human-LLM collaboration.
Finally, while our experiments rely on users reporting their trust levels in the AI assistant after each interaction, we also investigate whether these trust levels can be inferred solely from interaction features (Section 11). Our analysis of several heuristic and learning-based methods to estimate trust reveal that using purely interaction-level features is insufficient to accurately predict moments of low and high trust, thus highlighting the challenging nature of trust estimation and the need to consider other human factors.
In summary, our contributions are as follows:
  1. i)
    We first demonstrate that users’ trust level at the start of a given interaction affects reliance behavior (§5), with extreme levels of user trust resulting in worse decision-making.
  2. ii)
    Through controlled studies, we find that strategically providing supporting explanations when user trust is low and counter-explanations when trust is high reduces under-reliance and over-reliance, respectively (§6, 7, 8).
  3. iii)
    We also evaluate the utility of intervening by decelerating the interaction, finding that it helps reduce over-reliance but not under-reliance (§9).
  4. iv)
    We demonstrate that the findings from our controlled studies generalize when applied to a real Large Language Model (LLM) assisting users on decision-making tasks (§10).
  5. v)
    Finally, we explore several methods to estimate user trust based on interaction-level features, and find that these trust heuristics and models are insufficient for accurately predicting moments of low and high user trust, presenting open challenges in user trust modeling (§11).
@@
Our findings highlight the utility of modeling and adapting to user trust in AI-assisted decision-making, and present exciting avenues for improving human-AI collaboration.
2. Related Work
@@
We draw from a vast literature on measuring user trust in AI systems and evaluate how decision aids can be used for mitigating inappropriate reliance.
@@
Much work has explored the nature of human trust in AI systems (Lai et al., 2023), particularly in situations characterized by risk and uncertainty (Jacovi et al., 2021). Lee and See (2004) provide the most commonly accepted definition of trust, as a person’s attitude that an agent will help them achieve their goals. User trust is considered to be calibrated (Alizadeh et al., 2022) when it aligns with the AI system’s true capabilities (Wright, 2010), thus reducing AI misuse (e.g., over-reliance) and disuse (e.g., under-reliance) (Alizadeh et al., 2022). While trust in AI has typically been attributed to socio-economic and individual factors (Bach et al., 2024), recent work (Dhuliawala et al., 2023; Pareek et al., 2024) examines how trust develops as users interact with AI systems over multiple timesteps. Finally, while much work on trust in human-AI interaction is primarily concerned with trust in the AI’s ability to improve decision-making performance, trust could also concern other dimensions such as morality, fairness, and transparency (Malle and Ullman, 2021; Jacovi et al., 2021). In this work, we only explore the effect of performance-centric trust on decision-making.
@@
Bach et al. (2024) identify various mechanisms for measuring user trust, such as questionnaires (Schaffer et al., 2019), qualitative interviews (Barda et al., 2020), surveys (Lin et al., 2019), and point scales (Gulati et al., 2019). In AI-assisted decision-making, trust is often measured by observing user reliance behavior (Yin et al., 2019; Zhang et al., 2020); however, de Fine Licht and Brülde (2021) distinguish reliance, an observable behavior, from trust, a subjective belief.
Further, even when we are only concerned with performance-centric trust, user trust can be captured from multiple perspectives. The Trust in Automation (TiA) questionnaire (Körber, 2018) considers six aspects of user trust on system performance capabilities, including reliability/competence, predictability, user familiarity, perceptions of developers’ intentions, users’ propensity to trust, and overall trust in automation. In our user studies, we measure users’ perception of AI reliability on a 11-point scale, allowing users to report trust levels with more granularity than a traditional 5-point or 7-point Likert scale.
A final important design decision when measuring performance-centric trust is the granularity at which user trust is captured. Prior works have elicited local, interaction-specific trust levels by asking users to report their confidence in the AI’s accuracy for an individual question (Pareek et al., 2024; Dhuliawala et al., 2023). Others adopt a global lens by asking users to report their trust in the AI assistant’s general ability rather than in its performance for a specific problem (Körber, 2018; Schaffer et al., 2019). We elect to elicit global trust scores by asking users to report their belief in the AI’s helpfulness after each decision-making problem, and examine how this global trust varies and affects user behavior over the course of multiple interactions.
Figure 2. In our user study, each user interacts with an AI system for a sequence of $30$ decision-making problems. In each problem, the user first makes a decision by themselves, and then receives advice from the AI which they use to make a final decision. The user is then told what the correct decision is, and reports their trust in the AI (out of 10).
An illustration of our user study setup, where a user solves each decision-making problem in three phases.
@@
Inappropriate reliance, where users mistakenly accept incorrect AI predictions or reject correct ones (Parasuraman and Riley, 1997), is highly undesirable in high-stakes domains, such as healthcare and law (Schemmer et al., 2023). Appropriate reliance can be fostered through various decision aids, such as model confidences (Zhang et al., 2020), explanations (Wang and Yin, 2021), uncertainty expressions (Zhou et al., 2024; Kim et al., 2024), and providing sources (Feng and Boyd-Graber, 2019). Cognitive forcing functions (Buçinca et al., 2021), which insert friction (Chen and Schmidt, 2024; İnan et al., 2025) and promote deliberation (Park et al., 2019; Rastogi et al., 2022), are effective at mitigating over-reliance. We demonstrate how strategically providing these decision aids to users during moments of low or high trust can mitigate trust-induced inappropriate reliance.
Adaptive AI Assistance.
In the context of AI-assisted decision-making, most work on building adaptive AI assistants has focused on personalizing AI behavior to user attributes, such as their task capabilities (Ma et al., 2023), their need for cognition (Buçinca et al., 2024) and their tendency to over-rely (Swaroop et al., 2025). In all these works, the user attributes are known or determined before they interact with the AI assistant. In contrast, we adapt AI behavior to user trust levels, which develop and vary across time as the user interacts with the AI assistant over a sequence of decision-making problems.
Most similar to our work is Bhatt et al. (2025), which uses an online contextual bandit algorithm to decide which decision support policy to show to users, based on past interactions of the same user with different AI policies. This work rests on the assumption that the user’s reliance behavior on any given policy will stay stable across time. However, we argue that this assumption does not consider how past interactions with a policy will shape users’ trust in that policy, which in turn will alter future reliance behavior. Instead, we directly condition AI behavior on users’ current trust levels.
3. Preliminaries: Sequential AI-Assisted Decision-Making
In this work, we study how user trust impacts reliance on AI advice over a sequence of decision-making problems. We consider a setting where a human user interacts with an AI assistant on a sequence of $N$ decision-making problems from the same task, such as making medical diagnoses based on symptoms. Problems are tuples of input $x$, categorical choices $\mathcal{Y}$, and correct choice $y^{*}\in\mathcal{Y}$. The user solves each problem in three stages: independent decision-making based on their own knowledge, decision revision after viewing the AI advice, and trust update in the AI after observing decision accuracy (Figure 2). [^1] Our setup assumes that both the user and AI can observe the ground-truth decision after each problem.
@@ 1. Independent Decision-Making:
For the $i^{th}$ decision-making problem with input $x_{i}$, user makes a decision $y^{u,\text{init}}_{i}\in\mathcal{Y}$.
@@ 2. AI-Informed Decision Revision:
The user views the AI prediction $y^{AI}_{i}$ and confidence $c^{AI}_{i}\in[50\%,100\%]$, and subsequently makes a final decision $y^{u,\text{fin}}_{i}\in\mathcal{Y}$.
@@ 3. Trust Update:
After the user makes their final decision $y^{u,\text{fin}}_{i}$, they are informed of the accuracy of their decision and the AI prediction $y^{AI}_{i}$. Observing this feedback may alter the user’s trust in the AI assistant ’s ability to help them make better decisions. For instance, the AI misleading the user into making a wrong decision is likely to decay trust. After viewing the correctness of their decision, users report how much they trust the AI to help them with the decision-making task based on all user-AI interactions so far as an integer between 0 and 10.
@@
Our trust operationalization captures the user’s evolving belief in the AI’s helpfulness, which is likely to influence how the user relies on AI advice in subsequent decision-making problems.
Figure 3. An example of our study interface, where a user solves a decision-making problem from the ARC task in three stages.
An example of our user study interface
4. Experiment Methodology
4.1. Decision-Making Tasks
We conduct user studies on two decision-making tasks: ARC and Diagnosis. Table 7 contains examples of decision-making problems from both tasks. Figure 3 shows an example of how these decision-making problems are presented to users in our user study interface.
ARC Task.
This task consists of answering science questions. Questions are sourced from the ARC dataset (Clark et al., 2018), which consists of more than 7000 grade-school science multiple-choice questions written for examinations. The authors manually reviewed questions from this dataset and selected questions that were challenging (i.e., the correct answer is not immediately obvious, and at least one option was not obviously incorrect) but still understandable (i.e., did not contain any scientific jargon that laypeople may not be familiar with). All questions in the original dataset had four options, but the authors only selected the correct answer and the most plausible incorrect option for the decision-making problem. The final filtered set consisted of 39 questions.
Diagnosis Task.
This task consists of diagnosing patients based on patient age, gender, and symptoms. Patient symptoms are sourced from the DDXPlus dataset (Fansi Tchango et al., 2022), which contains 1.3 million synthetic patients with a differential diagnosis. The symptoms consist of an arbitrary number of descriptive statements, such as “The patient feels pain in their right knee.” We filter down to questions with only 10–15 intake responses so that users do not need to spend a long time reading and understanding the patient case. To convert the task into a multiple-choice problem, we select the top three negative conditions from the differential diagnosis as the incorrect options. Our final set of problems includes 55 cases corresponding to eleven different conditions.
4.2. Simulated AI Assistance
@@
Our user studies use a simulated AI assistant that provides a recommendation $R_{i}=(y^{AI}_{i},c^{AI}_{i})$} for each decision-making problem. We experiment with two types of AI assistants that vary in the quality of their confidence estimates: one assistant whose confidence is perfectly calibrated (Guo et al., 2017) with prediction accuracy, and one where the AI’s confidences are generally higher than their true accuracy (i.e. the AI is overconfident).
@@
This sampling procedure generates AI predictions and confidence scores for each decision-making problem.
4.3. Participant Recruitment
Participants were recruited on the Prolific platform. We recruited participants from the U.K. and U.S.A. who self-identified as fluent in English, and had at least 99% approval rate on previous studies.